6) Bad Company has to pick between two different machines that perform the same function, but have different lifespans. Machi I NEED A FULL STEP BY STEP PROCESS TO COMPLETION USING THE FORMULAS FROM "equivalent annual annuity chapter"
Answer #1

a) Initial cost =20000
Number of Years =3
Rate =10%
Operating costs per year =5000
PV of costs =Initial Cost +Operating Costs*((1-(1+r)^-n)/r =20000+5000*((1-(1+10%)^-3)/10% =32434.2599

b) Initial cost =15000
Number of Years =2
Rate =10%
Operating costs per year =7000
PV of costs =Initial Cost +Operating Costs*((1-(1+r)^-n)/r =15000+7000*((1-(1+10%)^-2)/10% =27148.7603

C)Equivalent annual annuity of A =PV of Costs of A/((1-(1+r)^-n)/r) =32434.2599/((1-(1+10%)^-3)/10%)=13042.30

d)Equivalent annual annuity of B =PV of Costs of A/((1-(1+r)^-n)/r) =27148.7603/((1-(1+10%)^-2)/10%)=15642.86

e)Machine A should be chosen because its has low equivalent annual cost